OpenStudy (anonymous):

The angle between 0 degrees and 360 degrees that is coterminal with the -1682 degrees angle is?

OpenStudy (anonymous):

Let the coterminal angle with -1682 degrees is x degrees therefore x-(-1682)= 360 x+1682=360 x= 360-1682=-1322